# Task Assignment and Exploration Optimization for Low Altitude UAV Rescue via Generative AI Enhanced Multi-agent Reinforcement Learning

### v1

@@ Abstract
The integration of emerging uncrewed aerial vehicle (UAVs) with artificial intelligence (AI) and ground-embedded robots (GERs) has transformed emergency rescue operations in unknown environments. However, the high computational demands of such missions often exceed the capacity of a single UAV, making it difficult for the system to continuously and stably provide high-level services. To address these challenges, this paper proposes a novel cooperation framework involving UAVs, GERs, and airships. This framework enables resource pooling through UAV-to-GER (U2G) and UAV-to-airship (U2A) communications, providing computing services for UAV offloaded tasks. Specifically, we formulate the multi-objective optimization problem of task assignment and exploration optimization in UAVs as a dynamic long-term optimization problem. Our objective is to minimize task completion time and energy consumption while ensuring system stability over time. To achieve this, we first employ the Lyapunov optimization technique to transform the original problem, with stability constraints, into a per-slot deterministic problem. We then propose an algorithm named HG-MADDPG, which combines the Hungarian algorithm with a generative diffusion model (GDM)-based multi-agent deep deterministic policy gradient (MADDPG) approach, to jointly optimize exploration and task assignment decisions. In HG-MADDPG, we first introduce the Hungarian algorithm as a method for exploration area selection, enhancing UAV efficiency in interacting with the environment. We then innovatively integrate the GDM and multi-agent deep deterministic policy gradient (MADDPG) to optimize task assignment decisions, such as task offloading and resource allocation. Simulation results demonstrate the effectiveness of the proposed approach, with significant improvements in task offloading efficiency, latency reduction, and system stability compared to baseline methods.
@@ 1 Introduction
Information collection and target detection from the disaster area is crucial to assess the situation and develop a suitable rescue plan. When ground transportation is interrupted by a disaster, it becomes difficult for human rescuers to enter the affected area. The existing method for rescuing in disaster areas involves using either ground vehicles or aerial robots that operate independently to gather data and transmit rescue information back to a ground commander [1]. Furthermore, ground-embedded robots (GERs) are used to explore for life forms and for environmental monitoring, while low-altitude uncrewed aerial vehicles (UAVs) play a significant role in post-disaster search and rescue activities. Emerging UAVs, combined with artificial intelligence (AI) technology, have revolutionized the way emergency rescue is handled [2, 3]. The UAVs offer several advantages in emergency rescue and disaster response. First, they can quickly access areas that are otherwise unreachable due to collapsed infrastructure or hazardous conditions, enabling rapid situational assessment. Second, UAVs equipped with high-resolution cameras and sensors can provide real-time aerial imagery and environmental data, greatly enhancing decision-making efficiency for rescue teams. Third, their flexibility in deployment allows for coordinated missions across wide areas, improving coverage and responsiveness. These capabilities make UAVs a valuable asset in time-sensitive and high-risk rescue operations. Despite the advantages of low-altitude UAVs for emergency rescue, there are still substantial challenges that need to be addressed.
@@
  • Solution: To address the complex optimization problem described above, this paper models the MOO problem as a markov decision process (MDP) and proposes a method based on the Hungarian algorithm and GDM-based multiagent deep deterministic policy gradient (MADDPG), named HG-MADDPG. Specifically, we introduce a novel application of the Hungarian algorithm for exploration area selection, which reduces the dimensionality of the observation space. Additionally, we integrate the GDM and MADDPG to enhance the generative decision-making capability of the actor network, enabling the optimization of task assignment.
@@
The rest of the paper is organized as follows. The related works are reviewed in section 2. In section 3, we present the system model and optimization problem. Section 4 reformulates the optimization problem. Section 5 introduces the novel method based on the HG-MADDPG method for task assignment and exploration optimization. Section 6 is dedicated to the simulation experiments and their analysis, while section 7 provides the concluding remarks.
@@ 2.2 MOO of MARL-based for UAV Computation Offloading
Task offloading in UAV networks using MARL is mainly divided into policy-based and value-based approaches. Examples of policy-based methods include MADDPG [27], multiagent twin delayed DDPG (MATD3)[28], and multiagent proximal policy optimization (MAPPO) [17]. The authors in [28] proposed a cooperative multi-agent deep reinforcement learning framework to derive the joint strategy for trajectory design, task allocation, and power management. In [27], the authors proposed a framework for multiaccess MEC in air–ground networks and employed a MADDPG algorithm for efficient, adaptive decision-making under complex constraints. The authors [17] proposed a UAV-assisted MEC network with a digital twin to enhance service for mobile users. They formulated a resource scheduling problem as an MDP and emphasized the role of MAPPO in optimizing computation offloading. The value-based methods primarily include QMIX [29] and value-decomposition networks (VDN) [21]. The authors [29] proposed aerial edge computing networks using UAVs for computation offloading. A MARL algorithm based on QMIX was introduced to manage the complexity of joint computation offloading and trajectory optimization. The authors [21] leveraged a MARL algorithm with value decomposition using a double deep Q-Network to optimize data aggregation and enhance offloading efficiency for UAV-enabled IoT systems in post-disaster contexts.
@@ 2.3 GAI-assisted Reinforcement Learning
Generative AI is a prominent subfield of machine learning focused on the conceptualization and creation of content. It arises from the goal of enabling machines to generate novel and original data that accurately reflects the underlying patterns, structures, and nuances present in the training datasets. Generative AI encompasses a variety of advanced models, including generative adversarial networks (GANs), transformer-based models, and GDM, each employing distinct methodologies for learning and generating data [30]. In [31], a GAN-driven auxiliary training mechanism for MARL is proposed. This approach reduces the overhead of real-world interactions and enables the agent’s policy to be well-suited for real-world execution environments by performing offline training using generated environment status. The authors in [32] combined Transformer with DRL to address the scalability of the network and proposed a Transformer-based MARL algorithm for scalable multi-UAV area coverage. In [33], the authors integrated AI-generated optimal decisions with DRL to develop the Deep Diffusion Soft Actor-Critic algorithm, improving the efficiency and effectiveness of selecting AI-generated content service providers.
@@ 3 System Model
Figure 1: System model. First, the GER computing power distribution map of the rescue area is divided into multiple subareas. Then, each UAV is assigned to multiple subareas. In each subarea, the UAV selects the most suitable GER based on real-time observations to approach and perform computing task offloading.
@@ 3.2 Mobility Model
We divide the system timeline $T$ into $I$ discrete time intervals, =$\mathcal{I}=\{1,2,\cdots,I\}$ , and $i\in\mathcal{I}$, each with equal length $\delta$, i.e., $T=I\delta$. Within the coverage area of an airship $g$, the numbers of UAVs and GERs are fixed for every time interval. Let $t_{0}$ signify the starting of $T$, and $t_{i}$ denote the time within the $i$-th interval, where $t_{i}\in[t_{0}+(i-1)\delta,t_{0}+i\delta]$. In the $i$-th time slot, the coordinate of UAV $u$ is given by $L_{u}(t_{i})=\left[x_{u}(t_{i}),y_{u}(t_{i}),H\right]$, where $[x_{u}(t_{i}),y_{u}(t_{i})]$ is the horizontal coordinate of UAV $u$ in the $i$-th time slot. The starting and ending coordinates of UAV $u$ are predefined, denoted as $L_{u}^{start}=\begin{bmatrix}x_{u}^{start},y_{u}^{start},H\end{bmatrix}$ and $L_{u}^{end}=\begin{bmatrix}x_{u}^{end},y_{u}^{end},H\end{bmatrix}$, respectively. The coordinates of UAV $u$ in the $i$-th time slot remain fixed if $\delta$ is sufficiently small. By combining the coordinates of UAV $u$ across $i$ time slots, the Euclidean distance between the starting and ending coordinates of UAV $u$ is as follows:
@@
The flight trajectory of UAV $u$ over time $T$ can be modeled as: $L_{u}=\left\{L_{u}^{start},L_{u}(t_{i}),L_{u}^{end}\right\},$ where $L_{u}(t_{i})=\left\{L_{u}(t_{1}),L_{u}(t_{2}),\cdots,L_{u}(t_{I})\right\}.$ The UAV’s flight trajectory is modeled as a set of flight segments corresponding to each time slot. In the $i$-th time slot, the status of UAV $u$ is expressed as:
@@
$\mathcal{L}_{t}$ The loss function for the denoising network model
@@ 3.3 Communication Model
$l_{u,j}^{{L}}(t_{i})=|\mathcal{H}_{u,j}^{{L}}(t_{i})|^{2}\left(\mathcal{I}_{u,j}^{{L}}(t_{i})\right)^{-1}10^{\frac{-g_{\sigma}^{{L}}(t_{i})}{10}},$ (4a)
$l_{u,j}^{{NL}}(t_{i})=|\mathcal{H}_{u,j}^{{NL}}(t_{i})|^{2}\left(\mathcal{I}_{u,j}^{{NL}}(t_{i})\right)^{-1}10^{\frac{-g_{\sigma}^{{NL}}(t_{i})}{10}},$ (4b)
where $\mathcal{H}_{u,j}^{L}(t_{i})$, $\mathcal{H}_{u,j}^{NL}(t_{i})$, $\mathcal{I}_{u,j}^{L}(t_{i})$, $\mathcal{I}_{u,j}^{NL}(t_{i})$, $g_{\sigma}^{L}(t_{i})$ and $g_{\sigma}^{NL}(t_{i})$ are the components of path loss, shadowing, and small-scale fading for LoS and NLoS links, respectively.
@@
$\mathcal{J}_{\sigma}^{{L}}(t_{i})\sim\mathcal{O}\left(0,\left(\sigma^{{L}}\right)^{2}\right),$ (7a)
$\mathcal{J}_{\sigma}^{{NL}}(t_{i})\sim\mathcal{O}\left(0,\left(\sigma^{{NL}}\right)^{2}\right),$ (7b)
@@ 3.5 Energy Consumption Model
Detection energy consumption. Throughout the entire flight of UAV $u$, there are unavoidable risks such as obstacles, adverse weather conditions, and complicated terrain. To estimate the energy expenditure for detecting risks during UAV’s flight, we define the risk sources as a set $\mathcal{Z}=\{1,2,\cdots,Z\}$, and $z\in\mathcal{Z}$. In this case, the UAV’s flight distance in the ${i}$-th time slot is divided into $m$ equal-length segments, and the Euclidean distance from the risk source $z$ to the central point of each segment is denoted by $\{d_{z_{1}},d_{z_{2}},\cdots,d_{z_{m}}\}$. When UAV $u$ travels from its starting coordinate $L_{u}^{start}$ to its ending coordinate $L_{u}^{end}$, its trajectory must be free of collisions. As such, UAV $u$ must monitor the specific rescue area of the risk source $z$ within each segment to ensure safe flying. Hence, the energy consumed by UAV $u$ in detecting the risk source $z$ during the ${i}$-th time slot is defined as $E_{u}^{z}(t_{i})=\varpi\sum_{k=1}^{m}d_{z_{k}}$, where $\varpi$ is the unit energy consumed by UAV $u$ for sensing the risk source $z$ [42]. When UAV $u$ encounters multiple risk sources from the set $\mathbb{Z}$ during flight, the total energy consumption for detection due to the surrounding risks can be expressed as:
@@ 4.2 Decoupling of P1 Via Lyapunov Algorithm
Algorithm 1 Lyapunov-Based Virtual Queue Online Control Algorithm
@@ 5.1 Overview of the HG-MADDPG Algorithm
The framework of the HG-MADDPG is shown in Fig. 3. The observation mechanism for the environment incorporates an area selection method based on the Hungarian algorithm, as illustrated in Algorithm 2. This method eliminates the need for the agent (i.e., UAV) to acquire the status of all GERs or the task data size across different areas, thereby conserving the computational resources of the UAV. The reverse diffusion process seeks to recover the original data from noisy observations. Specifically, its training process iteratively predicts the noise distribution and trains the reverse diffusion model, enabling the agent to extract more information about data distributions from the environment’s observations, generate actions, and execute them. The agent adjusts its RL parameters (such as denoising steps, batch size, and learning rate) based on feedback rewards, aiming to maximize long-term rewards and make optimal decisions. The task assignment and exploration optimization based on HG-MADDPG is detailed in Algorithm 3.
@@ 5.2.1 Motivation of Adopting Hungarian Algorithm
In low-altitude UAV systems with unknown environments, assigning flight areas effectively is crucial to maximize coverage and minimize redundancy or gaps. As the complexity of the environment increases, the observation space of agents expands, which introduces challenges for existing MARL methods like MADDPG. The motivation is driven by the need for optimal coverage while ensuring efficient UAV exploration. The Hungarian algorithm, with its optimal matching capability, ensures that each UAV is assigned to a specific area in a way that minimizes the total travel cost while covering all areas. Moreover, the Hungarian algorithm benefits from its low computational complexity, high efficiency, and stability, making it suitable for a variety of scenarios for solving matching problems [46]. Furthermore, the complexity is $O(\max(U,B)^{3})$, where $U$ is the number of UAVs and $\mathcal{B}$ is the number of subareas. This results in improved coverage efficiency, balanced workload distribution among UAVs, and reduced operational costs, making it a suitable choice for UAV exploration in unknown environments.

### v2

@@ Abstract
The integration of emerging uncrewed aerial vehicle (UAV) with artificial intelligence (AI) and ground-embedded robots (GERs) has transformed emergency rescue operations in unknown environments. However, the high computational demands of such missions often exceed the capacity of a single UAV, making it difficult for the system to continuously and stably provide high-level services. To address these challenges, this paper proposes a novel cooperation framework involving UAVs, GERs, and airships. This framework enables resource pooling through UAV-to-GER (U2G) and UAV-to-airship (U2A) communications, providing computing services for UAV offloaded tasks. Specifically, we formulate the multi-objective optimization problem of task assignment and exploration optimization in UAVs as a dynamic long-term optimization problem. Our objective is to minimize task completion time and energy consumption while ensuring system stability over time. To achieve this, we first employ the Lyapunov optimization method to transform the original problem, with stability constraints, into a per-slot deterministic problem. We then propose an algorithm named HG-MADDPG, which combines the Hungarian algorithm with a generative diffusion model (GDM)-based multi-agent deep deterministic policy gradient (MADDPG) approach, to jointly optimize exploration and task assignment decisions. In HG-MADDPG, we first introduce the Hungarian algorithm as a method for exploration area selection, enhancing UAV efficiency in interacting with the environment. We then innovatively integrate the GDM and multi-agent deep deterministic policy gradient (MADDPG) to optimize task assignment decisions, such as task offloading and resource allocation. Simulation results demonstrate the effectiveness of the proposed approach, with significant improvements in task offloading efficiency, latency reduction, and system stability compared to baseline methods.
@@ 1 Introduction
Information collection and target detection from the disaster area is crucial to assess the situation and develop a suitable rescue plan [1]. When ground transportation is interrupted by a disaster, it becomes difficult for human rescuers to enter the affected area. The existing method for rescuing in disaster areas involves using either ground vehicles or aerial robots that operate independently to gather data and transmit rescue information back to a ground commander [2, 3]. Furthermore, ground-embedded robots (GERs) are used to explore for life forms and for environmental monitoring, while low-altitude uncrewed aerial vehicles (UAVs) play a significant role in post-disaster search and rescue activities. Emerging UAVs, combined with artificial intelligence (AI), have revolutionized the way emergency rescue is handled [4, 5]. The UAVs offer several advantages in emergency rescue and disaster response. First, they can quickly access areas that are otherwise unreachable due to collapsed infrastructure or hazardous conditions, enabling rapid situational assessment. Second, UAVs equipped with high-resolution cameras and sensors can provide real-time aerial imagery and environmental data, greatly enhancing decision-making efficiency for rescue teams. Third, their flexibility in deployment allows for coordinated missions across wide areas, improving coverage and responsiveness. These capabilities make UAVs a valuable asset in time-sensitive and high-risk rescue operations. Despite the advantages of low-altitude UAVs for emergency rescue, there are still substantial challenges that need to be addressed.
@@
  • Solution: To address the complex optimization problem described above, this paper models the MOO problem as a markov decision process (MDP) and proposes a method based on the Hungarian algorithm and GDM-based multiagent deep deterministic policy gradient (MADDPG), named HG-MADDPG. Specifically, we introduce a novel application of the Hungarian algorithm for exploration area selection, which reduces the dimensionality of the observation space. The exploration process involves trajectory generation and obstacle detection. Additionally, we integrate the GDM and MADDPG to enhance the generative decision-making capability of the actor network, enabling the optimization of task assignment.
@@
The rest of the paper is organized as follows. The related works are reviewed in section 2. In section 3, we present the system model and optimization problem. Section 4 reformulates the optimization problem. Section 5 introduces the novel method based on the HG-MADDPG algorithm for task assignment and exploration optimization. Section 6 is dedicated to the simulation experiments and their analysis, while section 7 provides the concluding remarks.
@@ 2.2 MOO of MARL-based for UAV Computation Offloading
Task offloading in UAV networks using MARL is mainly divided into policy-based and value-based approaches. Examples of policy-based methods include MADDPG [29], multiagent twin delayed DDPG (MATD3)[30], and multiagent proximal policy optimization (MAPPO) [19]. The authors in [30] proposed a cooperative multi-agent deep reinforcement learning framework to derive the joint strategy for trajectory design, task allocation, and power management. In [29], the authors proposed a framework for multiaccess MEC in air–ground networks and employed a MADDPG algorithm for efficient, adaptive decision-making under complex constraints. The authors [19] proposed a UAV-assisted MEC network with a digital twin to enhance service for mobile users. They formulated a resource scheduling problem as an MDP and emphasized the role of MAPPO in optimizing computation offloading. The value-based methods primarily include QMIX [31] and value-decomposition networks [23]. The authors [31] proposed aerial edge computing networks using UAVs for computation offloading. A MARL algorithm based on QMIX was introduced to manage the complexity of joint computation offloading and trajectory optimization. The authors [23] leveraged a MARL algorithm with value decomposition using a double deep Q-Network to optimize data aggregation and enhance offloading efficiency for UAV-enabled IoT systems in post-disaster contexts.
@@ 2.3 GAI-assisted Reinforcement Learning
Generative AI is a prominent subfield of machine learning focused on the conceptualization and creation of content. It arises from the goal of enabling machines to generate novel and original data that accurately reflects the underlying patterns, structures, and nuances present in the training datasets. Generative AI encompasses a variety of advanced models, including generative adversarial networks (GANs), transformer-based models, and GDM, each employing distinct methodologies for learning and generating data [32]. In [33], a GAN-driven auxiliary training mechanism for MARL is proposed. This approach reduces the overhead of real-world interactions and enables the agent’s policy to be well-suited for real-world execution environments by performing offline training using generated environment status. The authors in [34] combined Transformer with deep reinforcement learning (DRL) to address the scalability of the network and proposed a Transformer-based MARL algorithm for scalable multi-UAV area coverage. In [35], the authors integrated AI-generated optimal decisions with DRL to develop the deep diffusion soft actor-critic algorithm, improving the efficiency and effectiveness of selecting AI-generated content service providers.
@@ 3 System Model
Figure 1: System model. First, the GER computing power distribution map of the rescue area is divided into multiple subareas. Then, each UAV is assigned to multiple subareas. In each subarea, the UAV selects a suitable GER based on real-time observations to approach and perform computing task offloading.
@@ 3.2 Mobility Model
We divide the system timeline $T$ into $I$ discrete time intervals, =$\mathcal{I}=\{1,2,\ldots,I\}$ , and $i\in\mathcal{I}$, each with equal length $\delta$, i.e., $T=I\delta$. Within the coverage area of an airship $g$, the numbers of UAVs and GERs are fixed for every time interval. Let $t_{0}$ signify the starting of $T$, and $t_{i}$ denote the time within the $i$-th interval, where $t_{i}\in[t_{0}+(i-1)\delta,t_{0}+i\delta]$. In the $i$-th time slot, the coordinate of UAV $u$ is given by $L_{u}(t_{i})=\left[x_{u}(t_{i}),y_{u}(t_{i}),H\right]$, where $[x_{u}(t_{i}),y_{u}(t_{i})]$ is the horizontal coordinate of UAV $u$ in the $i$-th time slot. The starting and ending coordinates of UAV $u$ are predefined, denoted as $L_{u}^{start}=\begin{bmatrix}x_{u}^{start},y_{u}^{start},H\end{bmatrix}$ and $L_{u}^{end}=\begin{bmatrix}x_{u}^{end},y_{u}^{end},H\end{bmatrix}$, respectively. The coordinates of UAV $u$ in the $i$-th time slot remain fixed if $\delta$ is sufficiently small. By combining the coordinates of UAV $u$ across $i$ time slots, the Euclidean distance between the starting and ending coordinates of UAV $u$ is as follows:
@@
The flight trajectory of UAV $u$ over time $T$ can be modeled as: $L_{u}=\left\{L_{u}^{start},L_{u}(t_{i}),L_{u}^{end}\right\},$ where $L_{u}(t_{i})=\left\{L_{u}(t_{1}),L_{u}(t_{2}),\ldots,L_{u}(t_{I})\right\}.$ The UAV’s flight trajectory is modeled as a set of flight segments corresponding to each time slot. In the $i$-th time slot, the status of UAV $u$ is expressed as:
@@
$\mathcal{L}_{t}$ The loss function for the denoising network
@@ 3.3 Communication Model
$l_{u,j}^{{L}}(t_{i})=|\mathcal{H}_{u,j}^{{L}}(t_{i})|^{2}\left(\mathcal{L}_{u,j}^{{L}}(t_{i})\right)^{-1}10^{\frac{-\mathcal{F}_{\sigma}^{{L}}(t_{i})}{10}},$ (4a)
$l_{u,j}^{{NL}}(t_{i})=|\mathcal{H}_{u,j}^{{NL}}(t_{i})|^{2}\left(\mathcal{L}_{u,j}^{{NL}}(t_{i})\right)^{-1}10^{\frac{-\mathcal{F}_{\sigma}^{{NL}}(t_{i})}{10}},$ (4b)
where $\mathcal{H}_{u,j}^{L}(t_{i})$, $\mathcal{H}_{u,j}^{NL}(t_{i})$, $\mathcal{L}_{u,j}^{L}(t_{i})$, $\mathcal{L}_{u,j}^{NL}(t_{i})$, $\mathcal{F}_{\sigma}^{L}(t_{i})$ and $\mathcal{F}_{\sigma}^{NL}(t_{i})$ are the components of path loss, shadowing, and small-scale fading for LoS and NLoS links, respectively.
@@
$\mathcal{F}_{\sigma}^{{L}}(t_{i})\sim\mathcal{O}\left(0,\left(\sigma^{{L}}\right)^{2}\right),$ (7a)
$\mathcal{F}_{\sigma}^{{NL}}(t_{i})\sim\mathcal{O}\left(0,\left(\sigma^{{NL}}\right)^{2}\right),$ (7b)
@@ 3.5 Energy Consumption Model
Detection energy consumption. Throughout the entire flight of UAV $u$, there are unavoidable risks such as obstacles, adverse weather conditions, and complicated terrain. To estimate the energy expenditure for detecting risks during UAV’s flight, we define the risk sources as a set $\mathcal{Z}=\{1,2,\ldots,Z\}$, and $z\in\mathcal{Z}$. In this case, the UAV’s flight distance in the ${i}$-th time slot is divided into $m$ equal-length segments, and the Euclidean distance from the risk source $z$ to the central point of each segment is denoted by $\{d_{z_{1}},d_{z_{2}},\ldots,d_{z_{m}}\}$. When UAV $u$ travels from its starting coordinate $L_{u}^{start}$ to its ending coordinate $L_{u}^{end}$, its trajectory must be free of collisions. As such, UAV $u$ must monitor the specific rescue area of the risk source $z$ within each segment to ensure safe flying. Hence, the energy consumed by UAV $u$ in detecting the risk source $z$ during the ${i}$-th time slot is defined as $E_{u}^{z}(t_{i})=\varpi\sum_{k=1}^{m}d_{z_{k}}$, where $\varpi$ is the unit energy consumed by UAV $u$ for sensing the risk source $z$ [44]. When UAV $u$ encounters multiple risk sources from the set $\mathbb{Z}$ during flight, the total energy consumption for detection due to the surrounding risks can be expressed as:
@@ 4.2 Decoupling of P1 Via Lyapunov Algorithm
Algorithm 1 Lyapunov-Based Virtual Queue Online Control
@@ 5.1 Overview of the HG-MADDPG Algorithm
The framework of the HG-MADDPG is shown in Fig. 3. The observation mechanism for the environment incorporates an area selection method based on the Hungarian algorithm, as illustrated in Algorithm 2. This method eliminates the need for the agent (i.e., UAV) to acquire the status of all GERs or the task data size across different areas, thereby conserving the computational resources of the UAV. The reverse diffusion process seeks to recover the original data from noisy observations. Specifically, its training process iteratively predicts the noise distribution and trains the reverse diffusion model, enabling the agent to extract more information about data distributions from the environment’s observations, generate actions, and execute them. The agent adjusts its parameters (such as denoising steps, batch size, and learning rate) based on feedback rewards, aiming to maximize long-term rewards and make optimal decisions. The task assignment and exploration optimization based on HG-MADDPG is detailed in Algorithm 3.
@@ 5.2.1 Motivation of Adopting Hungarian Algorithm
In low-altitude UAV systems with unknown environments, assigning flight areas effectively is crucial to maximize coverage and minimize redundancy or gaps. As the complexity of the environment increases, the observation space of agents expands, which introduces challenges for existing MARL methods like MADDPG. The motivation is driven by the need for optimal coverage while ensuring efficient UAV exploration. The Hungarian algorithm, with its optimal matching capability, ensures that each UAV is assigned to a specific area in a way that minimizes the total travel cost while covering all areas. Moreover, the Hungarian algorithm benefits from its low computational complexity, high efficiency, and stability, making it suitable for a variety of scenarios for solving matching problems [48]. Furthermore, the complexity is $O(\max(U,\mathcal{B})^{3})$, where $U$ is the number of UAVs and $\mathcal{B}$ is the number of subareas. This results in improved coverage efficiency, balanced workload distribution among UAVs, and reduced operational costs, making it a suitable choice for UAV exploration in unknown environments.